# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 2.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electric forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electric field (which can be of electrical or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electric field fluctuates, so their state is dynamic: the molecules dissociate, and the free ions recombine. Furthermore, the process plays an important role also in the energy supply of the neurons: the Adenozine Triphosphate (ATP) hydrolyzes molecules and the electric field near the membrane moves the ions to the "condenser plates". Depending on their environment, the substances can exist in base, ionized, and polarized states (and behaviors).
In biological cells, in a resting state, a small (approximately $10^{-3}$) portion of the molecules dissociates (i.e., the ions separate from their counterparts with opposite charges), and they move freely within the volume. In other words, the electrolyte liquid can then conduct electricity due to the mobility of the positive and negative ions (of course, if some electric/thermodynamic field provides the required driving force). The rest of the molecules can be more or less polarized, allowing for the possibility of producing an internal electric field in the solution. In a transition state, all components (including ions, molecules, and ATP) have gradients and they move towards their balanced state.
2.3 Deriving the "thermal electric field"
In a segmented electrolyte, the electric charges are typically globally balanced; however, they may become unbalanced locally due to physical reasons. The Nernst-Planck electrodiffusion equation in a balanced state
@@ 2.3 Deriving the "thermal electric field"
describes the mutual dependence of the spatial gradients of the electric and thermodynamic fields on each other. In good textbooks (see, for example, [25], Eq (11.28)), its derivation is exhaustively detailed. In the equation, $z$ is the spatial variable across the direction of the changed invasion parameter, $R$ is the gas constant, $F$ is the Faraday’s constant, $T$ is the temperature, $q$ the valence of the ion, $V_{m}(z)$} the potential, and $C_{k}(z)$} the concentration of the chemical ion. The two sides are the derivatives of the equation
@@
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine additively quantities from the intracellular and extracellular sides, whether concentrations, or mobilities, or Nernst voltages, as it happens in the GHK equation (see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [8]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [26], the two differentials must change in the same direction given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (2) results in a negative potential difference).
@@
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, furthermore, that U=EdU=E*d, we can derive the "equivalent thermodynamical electric field" for the case of a permeable ion channel in the membrane
@@
It does not depend on the concentrations, but depends linearly on the temperature and reciprocally on the membrane thickness (given in $nm$); furthermore, it is valid separately for all $C_{k}$
@@
Figure 1 depicts the field’s dependence on the membrane’s width and how it compares to the charge-generated electric field. The figure illustrates the setpoint, the boundaries of the changes, and that the interesting phenomena happen within the cross-section of the shaded areas. The figure is an illustration, but the figures are not far from the true ones. For a 5nm5\ nm membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Eq. (18). (Notice that the membrane’s thickness may change during the action potential, see [27]: the rush-in extra charge compresses the membrane. Similarly, the concentration of the electrolyte in proximity to the membrane also changes.) Correspondingly,
@@
Figure 1: The thermodynamic "electric field" depends on the membrane’s thickness, see Eq.(4). The electric gradient depends on the charged layer’s concentration and thickness; see Eq.(18). The dots mark measured data shown in Table 1.
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electric field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, joint charge-generated electric fields exist, and they equal, per ion, the concentration-generated "electric fields", see section 3.1.1.
@@
$0=\underbrace{e_{el}*E_{Gap}^{Total}(c,\Delta z)}_{Electrical~force}+\underbrace{e_{el}*E_{thermal}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{k}}}(d)}_{Thermodynamical~force}\quad\underbrace{\bigl(+F_{ext}(z)\bigr)}_{Constraint}\quad\underbrace{\bigl(+F_{Transp}(z)\bigr)}_{Transport~force}$ (6)
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added an external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [28], with a lot of cited references.)
@@
That means when describing an ionic transfer process, we must not separate the electric current from the mass transfer (in other words, to discuss electrical or thermodynamic models): they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electric term is not. To be entirely balanced, the system must be balanced to all elements. This way, changing one concentration implicitly changes all other concentrations and the electric field.
@@
Notice the special role of the constraint force by the closed volume represented by the membrane’s wall. The thermodynamic contribution also means a pressure from the colliding particles, and the electric repulsion exerts a force on the membrane’s wall. Those disciplinary forces act simultaneously and inseparably.
2.4 Fallacies about biological electricity
When discussing balanced states, no transport occurs, so the transport force cancels, and the mobility has no role (when speaking about mobility provided by protein mechanisms, one should derive the corresponding contribution to Eq. (7)). The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions on the opposite side, and the membrane mechanically provides the needed counterforce. This way, no resultant force acts on the ions; the two gradients persist. When the ions can move freely between the segments, they will continue to move until they produce a concentration gradient (a thermodynamic force) for the given ion that counterbalances the electrical gradient (the electrical force), and then the transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq. (27)) in a viscous fluid.
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electric current in solids.
It was a colossal mistake to forget that biological currents comprise ions with their electrostatic forces. It led to fallacies, among others, including leakage current (see Section 2.6.4) and heat absorption (see Section 4.6). Their finite (Stokes-Einstein) speed, see section 2.9, due to the ions’ mass and the different conduction mechanism, explains the spatiotemporal behavior of living matter [29] (contrasted to Newton’s ’instant interaction’).
It was another colossal mistake to introduce equivalent circuits with constant-value voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason, and prevents understanding how the competition between thermodynamic and electric processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 2.9; this way attributing the change of an electric entity to the biological material. This assumption neglects that a driving force is needed to move ions (denies Newton’s laws of motion). Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
The classic textbook [30], appendix A, correctly warned (especially when working with concepts of an electric theory of neuronal operation): "An intuitive grasp of concepts of electricity and electrical circuits is helpful for understanding some of the basic theory in cellular neurophysiology. …Under ideal circumstances, the physical act of measuring a neurophysiological event would have no effect on the electrical signal of interest. Unfortunately, this is seldom the case in neurophysiology."
As we discuss in section 4.1, physiology does not realize that in resting and transient states, different physical processes take place and correspondingly, different circuit models shall be used. In the resting state, the low intensity ion current is "shunted" by the ion channels in the membrane’s wall; moreover, the "parallel RCRC" model works sufficiently well. Since the current is slow, it is not simply distributed, but only the "overflow" reaches the AIS. However, as the current intensity exceeds the ion transmission capacity of the "resting ion channels", the "transient ion channels" play an increasing role. Above the voltage threshold, the serial $RC$ circuit is the correct model. Until the neuron’s voltage reaches its threshold, a "parallel" and a "serial" resistor work simultaneously, in this way implementing a mixed model. Measuring the "input resistance" of neurons actually measures, in this sense, a mixture of the "input" and "output" resistance corresponding to the two models, given that the current is distributed between the one through the membrane and the other through the AIS; plus the measuring electrode represents one more drain.
At the beginning, it was clearly told [31], but forgotten: "Since it is quite generally believed that the depolarization of a nerve fiber membrane, during excitation and propagation, involves an increased permeability to ions there have been many attempts to detect and measure this change as an increase in the electrical conductivity. … In these cases, the measuring current was also the stimulating current, and it was not possible to analyze the changes satisfactorily." Moreover, during the passage of an impulse, a "charge wave" passes along the axon and the changed number of charge carriers in that cross-section, see Eq.(20), changes the current that creates the illusion that the conductance at the measured cross section changed. In our terminology, the stabilizing feedback voltage provides an external force, and it also contributes an external current. The feedback amplifier’s electron current (converted to ion current) is provided externally and is added to the closed system, without including the measurement device into the closed system. That interpretation of a closed system denies the fundamental symmetries of science (among others, conservation of charge, mass, impulse, and energy). Moreover, it significantly contributes to the claim that "laws of science cannot describe life".
All conservation laws are correct and valid, only "the construction is different" [22] and the "test method" is wrong. Not mentioning that under some conditions, the "system under test" also "produces" charge carriers by opening ion channels. The "construction" needs revisiting concepts of electricity for living matter instead of projecting concepts valid for inanimate matter only, to living matter. That wrong "grasp of concepts of electricity" suggests that Ohm’s Law is not valid for living systems. The correct statement is that Ohm’s Law is not valid for arbitrarily chosen currents and voltages. Without that revisiting, by saying that living systems show non-Ohmic behavior, biology claims that it applies good laws inappropriately to a good system. The consequences are far-reaching: physiology uses magic changing conductances in circuits with fixed-voltage batteries, that blocks not only understanding the thermodynamic and mechanical effects (and introducing the corresponding theories [32, 33, 34, 35, 36, 27, 37, 38, 6, 39, 40, 41]), but even consistently describing the electric phenomena. We are not speaking against using the concept of conductance, but rather against deriving its value incorrectly by including foreign current compositions, which obscures the actual electrical behavior of biology. In addition, by introducing current feedback (which, by definition, compensates for any gradient in the system), physiology applies an opposite-phase control unit against the neuron’s native control unit and concludes that, after compensation, there is no gradient in the system; in other words, that life exists without needing a driving force.
Physiology is not aware that, inside the neuron, there is a condenser with a voltage difference between its plates. Instead, it believes that a ‘leakage current’ flows through resistors distributed over the membrane. That current is special in that, among other things, it has neither a chemical composition nor charge carriers; furthermore, it can not only emit but also absorb heat. Neuroscience observes the thermodynamic and mechanical changes (for a comprehensive review, see [33]) but cannot connect them to the electrical phenomena. Lacking the required physical background, biology claims that physical laws cannot describe processes in living matter, without providing a plausible and self-consistent theoretical model.
@@ 2.5 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electric effects connected to membranes and nanopores [19] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
@@
$\sigma_{V}(c)=c*N_{A}*e_{el}=c*6.023*10^{20}*1.602*10^{-19}=96.4*c\quad\biggl[\frac{C}{mM*m^{3}}\biggr]$ (8)
@@
We interpret classical physics’ notions from the theory of ’continuous’ electricity, interpreted for ’infinite’ cases, for the finite world of biological objects and the quantized atomic world. Our idea is similar to the one used by Boltzmann for connecting the continuous medium to quantized particles. When ions are present in an electrolyte having concentration $c$ in a ’conducting layer’ of thickness $\Delta z$ on an isolating surface, they produce a field
@@
Figure 2: The electric field of a disc of charge can be found by superposing the point charge fields of infinitesimal charge elements. This field can be facilitated by summing the fields of charged rings. “From Hyperphysics by Rod Nave, Georgia State University". ©hyperphysics.phy-astr.gsu.edu/hbase/
@@ 2.6 Two segments
Now, we separate the volume into two segments and compose the segments from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electric field’s $z$ component of those plates in the points as shown in Fig. 2. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electric field is zero inside the conducting layer), so the electric field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into the ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [42]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electric fields. Although not explicitly stated, we consider electric double layers to be present near the membrane, and ions from the dispersion medium are adsorbed on the particle’s surface, depending on the chemical properties of the medium.
@@
Correspondingly, the theory non-equivocally explains the potential near the membrane surface; see the discussion and review in [43], mainly their Figure 1. That model considers a "symmetric" condenser, that is, both plates contain charges of the same sign. The Donnan function is a (discontinuous) step-like potential. Its continuous version introduces a mathematically smooth, non-physical function at the boundary of the charged layer. One must describe a physically discontinuous layer (electrolyte, ion layer on the membrane’s surface, and isolating membrane) with a single continuous function and gradients (potential and concentration) with physical meaning.
@@
The behavior near the boundary layers is problematic, primarily because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electric field; the charged layer and the electrolyte jointly produce the electric field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
@@
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electric field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their electric force is a fragment of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it generates an electric field similar to that of a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electric field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separates the segments). The membrane does not need to generate any counterforce (except for collisions).

### v3

@@ 4.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electrical forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electrical field (which can be of electrical, magnetic, mechanical, or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electrical field fluctuates, so their state is dynamic: molecules dissociate, and free ions recombine. Furthermore, the process also plays an important role in the energy supply of neurons: Adenozine Triphosphate (ATP) hydrolyzes molecules, and the electrical field near the membrane moves ions to the "condenser plates", thereby accumulating "potential energy". (The claim that the ions passing through ion channels do not use energy is wrong: they are accelerated by the condenser’s field and consume energy in that way.) Depending on their environment, substances can exist in base, ionized, and polarized states (and exhibit corresponding behaviors).
(For the sake of clarity, both mechanisms are known in physics. The term ’polarization’ is reserved for the redistribution of charges in multiply charged systems without the application of an external current. Changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even though, unfortunately, physiology refers to it as polarization. As discussed below, true polarization is also present, which confuses the description. External currents contribute to the changes of potential.)
In biological cells, in a resting state, a small (approximately $10^{-3}$) fraction of the molecules dissociates (i.e., ions separate from their counterparts with opposite charges) and moves freely within the volume. In other words, the electrolyte liquid can conduct electricity (under the effect of an electrical or thermodynamic field) because the positive and negative ions are mobile. The remaining molecules can be more or less polarized, allowing the possibility of generating an internal electrical field in the solution. In a transition state, all components (including ions, molecules, and ATP) have gradients, and they move towards their balanced state.
4.3 Deriving the "thermodynamic electrical field"
In a segmented electrolyte, electrical charges are typically globally balanced; however, they may become locally unbalanced due to physical factors. The Nernst-Planck electrodiffusion equation in a balanced state
@@ 4.3 Deriving the "thermodynamic electrical field"
describes the mutual dependence of the spatial gradients of the electrical and thermodynamic fields on each other. In good textbooks (see, for example, [34], Eq (11.28)), its derivation is exhaustively detailed. In the equation, $z$ is the spatial variable across the direction of the changed invasion parameter, $R$ is the gas constant, $F$ is the Faraday’s constant, $T$ is the temperature, $q$ the valence of the ion, $V_{m}(z)$} the potential, and $C_{k}(z)$} the concentration of the chemical ion. The two sides are the derivatives of the equation
@@
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine quantities from the intracellular and extracellular sides additively, whether concentrations, mobilities, or Nernst voltages, as it happens in the Goldman-Hodgkin-Katz (GHK) equation (see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [31]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [6], the two differentials must change in the same direction, given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (2) results in a negative potential difference). (The derivation in ’ordinary physics’, such as  [34], Eq.(11.30) is wrong, at least because of two reasons: the same speed is assumed for mass and charge transport; furthermore, the used partial derivatives are not independent; the details are discussed throughout the paper.)
@@
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, furthermore, that U=EdU=E*d, we can derive the "equivalent thermodynamic electrical field" for the case of a permeable ion channel in the membrane
@@
It does not depend on concentration, but rather is linear in temperature and inversely proportional to membrane thickness (given in $nm$). Furthermore, it is valid separately for all $C_{k}$
@@
Figure 1 depicts the field’s dependence on the membrane’s width and how it compares to the charge-generated electrical field. The figure illustrates the setpoint, the boundaries of the changes, and the interesting phenomena that occur within the shaded cross-section. The figure is an illustration, but its figures are not far from the true ones. For a 5nm5\ nm membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Eq. (18). (Notice that the membrane’s thickness may change during the action potential, see [35]: the rush-in extra charge compresses the membrane. Similarly, the concentration of the electrolyte in proximity to the membrane also changes.) Correspondingly,
@@
Figure 1: The thermodynamic "electrical field" depends on the membrane’s thickness; see Eq.(4). The electrical gradient depends on the charged layer’s concentration and thickness; see Eq.(18). The dots mark measured data shown in Table 1.
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electrical field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, joint charge-generated electrical fields exist, and they are equal, per ion, to the concentration-generated "electrical fields", see section 5.1.1.
@@
$0=\underbrace{e_{el}*E_{Gap}^{Total}(c,\Delta z)}_{Electrical~force}+\underbrace{e_{el}*E_{thermal}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{k}}}(d)}_{Thermodynamic~force}\quad\underbrace{\bigl(+F_{ext}(z)\bigr)}_{Constraint}\quad\underbrace{\bigl(+F_{Transp}(z)\bigr)}_{Transport~force}$ (6)
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added a transport-related external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [36], with a lot of cited references.) This equation, expressing Newton’s first law, has no place for effects such as "Pumps that maintain ion gradients …transport ions against their electrical and chemical gradients[17], page 101. It implies some new force that moves ions, confronting Schrödinger’s science-based opinion: "And that not on the ground that there is any ‘new force’ or whatnot, directing the behaviour of the single atoms within a living organism"[7]. Biophysics still did not provide either the origin of the ’new force’ or the reason why Newton’s laws are not valid for ions in living matter, which undermines the credibility of those protein-based mechanisms.
@@
That means when describing an ionic transfer process, we must not separate the electrical current from the mass transfer (in other words, to discuss electrical or thermodynamic models in a disciplinary way): they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electrical term is not. To be entirely balanced, the system must be balanced to all elements. This way, changing one concentration implicitly changes all other concentrations and the electrical field.
@@
Notice the special role of the constraint force by the closed volume represented by the membrane’s wall. The thermodynamic contribution also means a pressure from the colliding particles, and the electrical repulsion exerts a force on the membrane’s wall. Those disciplinary forces act simultaneously and inseparably.
@@ 4.4 One segment
For the discussion below, we assume that the segment has a two-dimensional surface boundary, and we discuss the gradients along a line perpendicular to that plane’s surface. We neglect the electrokinetic and surface-related electrical effects connected to membranes and nanopores [27] and focus on the balanced state of membranes in ionic solutions. We compose the segments from such layers (sheets), $(x,y)$ parallel plates, and describe the gradients in the direction of $z$. We introduce the concept of a ’thin physical layer’ parallel to the membrane, characterized by a finite width.
@@
$\sigma_{V}(c)=c*N_{A}*e_{el}=c*6.023*10^{20}*1.602*10^{-19}=96.5*c\quad\biggl[\frac{C}{mM*m^{3}}\biggr]$ (8)
@@
We interpreted classical physics’ notions from the theory of ’continuous’ electricity, interpreted for ’infinite’ cases, for the finite world of biological objects and the quantized atomic world. Our idea is similar to Boltzmann’s for connecting the continuous medium to particles. When ions are present in an electrolyte having concentration $c$ in a ’conducting layer’ of thickness $\Delta z$ on an isolating surface, they produce a field
@@
Figure 2: The electrical field of a disc of charge can be found by superposing the point charge fields of infinitesimal charge elements. This field can be facilitated by summing the fields of charged rings. “From Hyperphysics by Rod Nave, Georgia State University". ©hyperphysics.phy-astr.gsu.edu/hbase/
@@ 4.5 Two segments
Now, we split the volume into two segments and compose them from ’thin physical layers’ (sheets) with different potentials (and concentrations). We consider the immediate environment of the neuron’s membrane as adjacent parallel $(x,y)$ plates and find the electrical field’s $z$ component of those plates in the points as shown in Fig. 2. In classic electricity, the charge remains on the surface of the conducting layer, and the distribution is homogeneous (although the charge is concentrated on the surface and the electrical field is zero inside the conducting layer). Hence, the electrical field contributions cancel each other. In electrolyte segments, (a tiny fraction of) the molecules decompose into the ionized state (dissociate), and the created ions interact with a bounding membrane using a not entirely understood mechanism [37]. The rest of the dipole molecules become polarized but do not dissociate, forming virtual charges that allow the electrolytes to contain electrical fields. Although not explicitly stated, we assume that electrical double layers are present near the membrane, and ions from the dispersion medium are adsorbed onto the particle’s surface, depending on the medium’s chemical properties.
@@
Correspondingly, the theory non-equivocally explains the potential near the membrane surface; see the discussion and review in [38], mainly their Figure 1. That model considers a symmetrical condenser, that is, both plates contain charges of the same sign. The Donnan function is a (discontinuous) step-like potential. Its continuous version introduces a mathematically smooth, non-physical function at the boundary of the charged layer. One must describe a physically discontinuous layer (electrolyte, ion layer on the membrane’s surface, and isolating membrane) with a single continuous function and gradients (potential and concentration) with physical meaning.
@@
The behavior near the boundary layers is problematic, primarily because the layers themselves are ill-defined. Physically, they are not ideally plain and well-defined surfaces; the layer thicknesses are limited by the size of the ions/molecules and their associated ions; the layers do not necessarily satisfy the ideal requirements of the respective physical laws. Mathematically, a single ’universal’ function that spans all regions surely cannot accurately describe the behavior of the physical model. So, we chose a piecewise description where the pieces seamlessly fit at the boundaries. We assume that between the surfaces of the membrane, classical electricity describes the electrical field; the charged layer and the electrolyte jointly produce the electrical field; we have a uniformly charged layer from separated charges on the surface and a polarized electrolyte outside them.
@@
Unlike in the classical theory, the electrolyte segment next to the surface layer contains dipole molecules that have more or less balanced charges (the polarization also depends on the external electrical field), so they have much less mobility than the dissociated ions: their size and mass are bigger, and their charge is a fraction of that of the ions. Their virtual charge becomes apparent only in the appropriate environment, yet it generates an electrical field similar to that of a real charge. The final reasons for creating virtual charges are real charges and/or external fields, so an electrical field due to virtual charges can also exist inside dielectric matter. We assume the membrane is transparent for the electrical interaction (the electrical field affects the ions in the other segment on the other side of the membrane) but not for their masses (mechanically separate the segments). The membrane does not need to generate any counterforce (except for collisions).